OpenStudy (anonymous):

1.There are 10 kids running a race. How many ways are there for them to come in first, second, and third place? 2.In a drawing, a jar holds the numbers 1 to 8. If you draw three numbers to form a three-digit number (without replacement), what is the probability you will draw 135?

OpenStudy (amistre64):

10.9.8 = 720

OpenStudy (amistre64):

8! = 40320 i think

OpenStudy (amistre64):

1/8! perhaps? does the 135 have to be in that order?

OpenStudy (anonymous):

\[\frac{1}{8}\times \frac{1}{7}\times \frac{1}{6}\]

OpenStudy (amistre64):

ack .... satellite got it lol

OpenStudy (anonymous):

i am assuming you draw a 1, then a 3, then a 5 in that order, yes

OpenStudy (anonymous):

i get one in at sometimes.

OpenStudy (amistre64):

i drew 8 numbers ;)

OpenStudy (anonymous):

lol the rest were zeros!